find the sinking fund payment

6.) cavell manufacturing established a slinking fund to pay off a 100,000,000 bond issue that comes due in 8 years. A) what equal payments must be deposited in the fund every 3 months at 6% interest compounded quarterly in order for cavell to meet this financial obligation ? B) what is the total amount of interest earned in the sinking fund account?

7) the Carmichael hotel has a financial obligation of 10,000,000 due in 5 years. A sinking fund is established to meet this obligation at 12% interest compounded monthly .a) what equal monthly sinking fund payments are required to accumulate the needed amount? B.) what is the total amount of interest earned in the account?


1
Expert's answer
2021-05-19T17:50:06-0400

6a) Equal payment: bond issues ÷ years in month × rate

100000000÷(8×12×3) 6/100

(100000000÷17.28 )= $5787037

6b) Total amount of interest earned in the sinking fund account=

(100000000- 5787037)= $ 94212.963

7a) Equal monthly payment

Bond issues ÷ years in month × rate

10000000÷ (5×12) × 12/100

(10000000÷ 7.2)= $ 1388888

7b) Total amount of interest

(10000000- 1388888)= $ 8611112